# The Impact of Differential Feature Under-reporting on Algorithmic Fairness

### v1

@@ Results
Figure F.9 depicts a subset of the results for the COMPAS dataset. Comparing against the results of the noise-free setting summarized in Figure D.1, we observe that the group Other is under-selected to a greater extend with additional noise. Under-selection occurs even if no feature under-reporting is added (dark blue curves) and increases with increasing noise, i.e. decreasing $R^{2}$ of the model on $Z$. On a high level, this occurs because the predictions $\hat{Y}_{Z}$ concentrate more closely around their group-level means as compared to the true values $Y$. The mean of $\hat{Y}_{Z}$ is smaller for the group Other than the group African-American which leads to under-selection of the group Other as compared to the true $Y$ at many thresholds. We note that the group-level variances in outcomes $Y$ and predictions $\hat{Y}_{Z}$ play a role in this dynamic as well. The isolated effect of feature under-reporting in the studied setting appears to be similar to the effect in the noise-free setting. As under-reporting is introduced into the group Other via the feature ‘priors count’, the group Other is further under-selected. The more under-reporting is injected, the more the group is under-selected. The magnitude of under-selection due to under-reporting is comparable across different levels of regression noise. Overall, the results give us some insight into what to expect in more realistic settings of feature under-reporting. Instead of selection rate disparities that are exclusively due to differential feature under-reporting, disparities in the studied setting also depend on regression noise which, together, leads to increased disparities overall.
F.3. Possibility of decreasing disparities
@@
Figure F.10 depicts the parameter estimates and excess selection rate of group Other (i.e. not African-American) when fitting a model on data with feature under-reporting in the feature ‘age’ for group Other. We see that under-reporting in this setting leads to over-selection of the group with under-reporting. This over-selection is increasing with increasing levels of under-reporting. As the figure shows, the regression parameter for age is negative with an attenuation effect when under-reporting is injected. This means that in the semi-synthetic ground truth model and in the prediction models under under-reporting younger defendants are more likely to reoffend than older defendants. The feature correlations between age and juvenile crime counts (felony, misdemeanor, and other) are negative in the data while the correlation between age and the feature ‘priors count’ is positive. This leads to parameter estimates that are increasing for increasing under-reporting in age for juvenile crime counts and decreasing for increasing under-reporting for priors count exactly as predicted by the theoretical analyses in Proposition 3. Ultimately, this example shows how, in some settings, disparities may decrease as a function of under-reporting which aligns with Case 1 from the theoretical discussions in Section 5. However, the example presented here is somewhat artificial and we find that typically disparities are increasing with differential feature under-reporting.
Figure F.10. Excess selection rate of group Other (i.e. not African-American) at different population selection rates with synthetic outcomes using the COMPAS dataset, and the respective parameter estimates. Under-reporting is added to the feature ‘age’ in group Other. Results are reported as averages over 30 simulation runs with shaded areas representing one standard deviation in each direction. Note that parameter estimates are only displayed for continuous count features and age to preserve readability. The models additionally take sex and the categorical feature charge degree into consideration.

### v2

@@ Results
Figure G.9 depicts a subset of the results for the COMPAS dataset. Comparing against the results of the noise-free setting summarized in Figure E.1, we observe that the group Other is under-selected to a greater extend with additional noise. Under-selection occurs even if no feature under-reporting is added (dark blue curves) and increases with increasing noise, i.e. decreasing $R^{2}$ of the model on $Z$. On a high level, this occurs because the predictions $\hat{Y}_{Z}$ concentrate more closely around their group-level means as compared to the true values $Y$. The mean of $\hat{Y}_{Z}$ is smaller for the group Other than the group African-American which leads to under-selection of the group Other as compared to the true $Y$ at many thresholds. We note that the group-level variances in outcomes $Y$ and predictions $\hat{Y}_{Z}$ play a role in this dynamic as well. The isolated effect of feature under-reporting in the studied setting appears to be similar to the effect in the noise-free setting. As under-reporting is introduced into the group Other via the feature ‘priors count’, the group Other is further under-selected. The more under-reporting is injected, the more the group is under-selected. The magnitude of under-selection due to under-reporting is comparable across different levels of regression noise. Overall, the results give us some insight into what to expect in more realistic settings of feature under-reporting. Instead of selection rate disparities that are exclusively due to differential feature under-reporting, disparities in the studied setting also depend on regression noise which, together, leads to increased disparities overall.
G.3. Possibility of decreasing disparities
@@
Figure G.10 depicts the parameter estimates and excess selection rate of group Other (i.e. not African-American) when fitting a model on data with feature under-reporting in the feature ‘age’ for group Other. We see that under-reporting in this setting leads to over-selection of the group with under-reporting. This over-selection is increasing with increasing levels of under-reporting. As the figure shows, the regression parameter for age is negative with an attenuation effect when under-reporting is injected. This means that in the semi-synthetic ground truth model and in the prediction models under under-reporting younger defendants are more likely to reoffend than older defendants. The feature correlations between age and juvenile crime counts (felony, misdemeanor, and other) are negative in the data while the correlation between age and the feature ‘priors count’ is positive. This leads to parameter estimates that are increasing for increasing under-reporting in age for juvenile crime counts and decreasing for increasing under-reporting for priors count exactly as predicted by the theoretical analyses in Proposition 3. Ultimately, this example shows how, in some settings, disparities may decrease as a function of under-reporting which aligns with Case 1 from the theoretical discussions in Section 5. However, the example presented here is somewhat artificial and we find that typically disparities are increasing with differential feature under-reporting.
Figure G.10. Excess selection rate of group Other (i.e. not African-American) at different population selection rates with synthetic outcomes using the COMPAS dataset, and the respective parameter estimates. Under-reporting is added to the feature ‘age’ in group Other. Results are reported as averages over 30 simulation runs with shaded areas representing one standard deviation in each direction. Note that parameter estimates are only displayed for continuous count features and age to preserve readability. The models additionally take sex and the categorical feature charge degree into consideration.